McGilla Golf has decided to sell a new line of golf clubs. The clubs will sell for $750 per set and have a variable cost of $350 per set. The company has spent $145,000 for a marketing study that determined the company will sell 57,000 sets per year for seven years. The marketing study also determined that the company will lose sales of 9,000 sets of its high-priced clubs. The high-priced clubs sell at $1,050 and have variable costs of $650. The company will also increase sales of its cheap clubs by 10,500 sets. The cheap clubs sell for $390 and have variable costs of $205 per set. The fixed costs each year will be $9,050,000. The company has also spent $1,060,000 on research and development for the new clubs. The plant and equipment required will cost $28,350,000 and will be depreciated on a straight-line basis. The new clubs will also require an increase in net working capital of $1,250,000 that will be returned at the end of the project. The tax rate is 40 percent, and the cost of capital is 10 percent.

    

Suppose you feel that the values are accurate to within only ±10 percent. What are the best-case and worst-case NPVs? (Hint: The price and variable costs for the two existing sets of clubs are known with certainty; only the sales gained or lost are uncertain.) (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

   

                     NPV
  Best-case $   
  Worst-case $   
Answer #1
Costs of marketing study and Research are sunk costs
These costs are already incyrred and not relevant for the analysis
BEST CASE SCENARIO:
A Sales Price per set $825 (750*1.1)
B Variable Costs per set $315 (350*0.9)
C=A-B Contribution margin per set $510
D Quantity of sales per year                     62,700 (57000*1.1)
E=C*D Annual Contribution $31,977,000
F Contribution margin of high priced clubs $400 (1050-650)
G Quantity of lost sales of high priced                       8,100 (9000*0.9)
H=F*G Loss of contribution per annum $3,240,000
I Contribution margin of Cheap clubs $185 (390-205)
J Quantity of Increase in sales of Cheap clubs                     11,550 (10500*1.1)
K=J*I Addition to contribution per annum $2,136,750
L=E-H+K Total AnnualContributon $30,873,750
M Annual Fixed Costs $8,145,000 (9050000*0.9)
N=L-M Annual Profit (excluding depreciation) $22,728,750
P Cost of new equipment $        25,515,000 (28350000*0.9)
Q=P/7 Depreciation per year $          3,645,000
R=N-Q Before tax operating profit $19,083,750
S=R*(1-0.4) After tax operating profit(40% tax rate) $11,450,250
T=S+Q After tax annualoperating cash flow $15,095,250
PV1 Present Value of annual operating cash flow $73,489,999 (Using PV function of excelwith Rate=10%,Nper=7, Pmt=-T)
U Increase in net working capital $1,125,000 (1250000*0.9)
PV2 Present Value of returned working capital $              577,303 (1125000/(1.1^7)
NPV=PV1+PV2-P-U NET PRESENT VALUE (NPV) BEST CASE $47,427,302
WORST CASE SCENARIO:
A Sales Price per set $675 (750*0.9)
B Variable Costs per set $385 (350*1.1)
C=A-B Contribution margin per set $290
D Quantity of sales per year                     51,300 (57000*0.9)
E=C*D Annual Contribution $14,877,000
F Contribution margin of high priced clubs $400 (1050-650)
G Quantity of lost sales of high priced                       9,900 (9000*1.1)
H=F*G Loss of contribution per annum $3,960,000
I Contribution margin of Cheap clubs $185 (390-205)
J Quantity of Increase in sales of Cheap clubs                       9,450 (10500*0.9)
K=J*I Addition to contribution per annum $1,748,250
L=E-H+K Total AnnualContributon $12,665,250
M Annual Fixed Costs $9,955,000 (9050000*1.1)
N=L-M Annual Profit (excluding depreciation) $2,710,250
P Cost of new equipment $        31,185,000 (28350000*1.1)
Q=P/7 Depreciation per year $          4,455,000
R=N-Q Before tax operating profit ($1,744,750)
X Tax saving for loss $697,900
S=R+X After tax operating profit(40% tax rate) ($1,046,850)
T=S+Q After tax annualoperating cash flow $3,408,150
PV1 Present Value of annual operating cash flow $16,592,302 (Using PV function of excelwith Rate=10%,Nper=7, Pmt=-T)
U Increase in net working capital $1,375,000 (1250000*1.1)
PV2 Present Value of returned working capital $              705,592 (1375000/(1.1^7)
NPV=PV1+PV2-P-U NET PRESENT VALUE (NPV) WORST CASE ($15,262,106)
NPV
Best Case $47,427,302
Worst case ($15,262,106)